OpenStudy (is3535):

What is the exact circumference of a circle with a radius of 12 inches? A. 6 inches B. 12 inches C. 24 inches D. 44inches

OpenStudy (is3535):

@ElonaSushchik

OpenStudy (elonasushchik):

B. 12 inches

OpenStudy (is3535):

are you 100% sure

OpenStudy (elonasushchik):

evidence so can i have my cookies

OpenStudy (is3535):

yea

OpenStudy (anonymous):

the radius is 12 not the circumference

OpenStudy (is3535):

so not 12 @jrain35

OpenStudy (elonasushchik):

Diameter = 12 inches, Circumference = 12*(pi) = approx. 37.699 Radius = 12 inches, Circumference = 2*12*(pi) = 24*(pi) = approx. 75.398 inches If the author has actually measured the outside of the circle (i.e. the perimeter of a dinner plate), that is the circumference! Problem solved.

OpenStudy (anonymous):

diameter is 24 if radius is 12

OpenStudy (is3535):

so 24